# QG from SymQRG: AdS3/CFT2 Correspondence as Topological Symmetry-Preserving Quantum RG Flow

### v1

@@ 3 $T\bar{T}$ deformation for 2D CFTs as SymQRG
which is precisely the solution of (3.1) in terms of the Hubbard-Stratonovich transformation, and can be interpreted as coupling the original theory to 2D random geometry[85, *Cardy:2018sdv, 32].
@@ 3.3 Wheeler-DeWitt equation from $T\bar{T}$ deformation
with the integration limit switched to characterise evolution into the bulk. Then we introduce a collection of states $\bra{g_{ij}=\frac{\gamma_{ij}}{\lambda},\lambda}$ |, in analogous to the “position basis” in quantum mechanics, to be metric eigenstates in the Heisenberg picture, which plays the role of ${}_{\Lambda}\bra{\Omega^{\lambda}(\gamma_{ij})}$ | discussed in the earlier section. Then the Schrödinger picture metric eigenstates are defined at reference “time” $\lambda=\epsilon$ satisfying [68, *PhysRevD.28.2960]:
@@
The state $\bra{g_{ij}=\frac{\gamma_{ij}}{\mu},\mu}$ | corresponds to the standard Dirichlet boundary conditions imposed on finite-cutoff surfaces in the AdS/CFT correspondence. These conditions can equivalently be described as mixed boundary conditions at asymptotic infinity, linked to double-trace deformations [93, 72, 73, *Guica:2022gts]. Given that $T\bar{T}$-deformed theories have been demonstrated to preserve the Virasoro algebra in a highly nontrivial manner [73, *Guica:2022gts], it implies that these boundary conditions also lead to the same Virasoro symmetry.
@@ 4.2 SymQRG for tensor network state sum representation of 2D CFT path integrals
The observation made in [33, 34, 35, 1], motivated by [16, 100, 17] and the foundational works of FRS [36, *Fuchs:2003id, *Fuchs:2003id1, *Fuchs:2004xi, *Fjelstad:2005ua, *Fjelstad:2006aw], and also independently in [101, *Brehm:2024zun], is that 2D CFTs admit a discrete tensor network state sum representation arising from a novel perspective on cutting and gluing their path integrals. The result can be explicitly written as an overlap (2.5) in the Hilbert space of the Levin-Wen string net models.
@@
The tilde indicates a special choice of normalization that we explain in (B.3). We introduce the diagrammatic notation, where we denote $\tilde{\gamma}^{\alpha_{1}\alpha_{2}\alpha_{3}}_{I_{1}I_{2}I_{3}}(R)$ using black dashed lines, and $\tilde{C}^{\sigma_{3}\sigma_{1}\sigma_{2}}_{\alpha_{1}\alpha_{2}\alpha_{3}}$ using blue solid lines and red circles, as shown in Fig. 10.
For general (non-diagonal) CFTs, there is a remarkable formula first discovered by Runkel, relating the BCFT OPE coefficients to the (generalized) $6j$ symbols[103, *Runkel:1999dz, 36] as,
@@
As constructed, both states defined on the discrete lattice are projected entangled pair state (PEPS) tensor network states [^21] For readers unfamiliar with PEPS tensor networks, they can be understood as a two-dimensional generalization of the more familiar one-dimensional matrix product states(MPS) [108, *8589270]. defined on a two-dimensional lattice [110, *Verstraete:2004cf], as illustrated in Fig. 13. The dangling physical legs, shown in red, correspond to the primary labels $\alpha_{i}$. The hidden layer of the auxiliary legs represents the internal degrees of freedom, and they take values in the set of boundary conditions $\sigma_{a}$ for |$\ket{\Psi}_{\Lambda}$ and in the set of integer-valued descendant labels $I_{i}$ for ${}_{\Lambda}\bra{\Omega^{R}}$ |. For ${}_{\Lambda}\bra{\Omega^{R}}$ |, the presence of infinitely many descendants results in a bond dimension that is infinite. This infinite bond dimension is what allows for an exact representation of the fixed-point tensor network of the gapless boundary condition associated with a continuum field theory.
@@
The idea of rewriting a 2D field theory partition function as a “strange correlator”(2.5) in the topologically ordered string-net Hilbert space, to capture the topological symmetry of the theory, was inspired by the proposal in [16]. This approach was initially applied to certain 2D integrable lattice models, generalizing similar observations made for symmetry protected topological (SPT) phases in [112]. A key insight from [16] is that this formulation naturally provides an analytic real-space tensor network RG algorithm, taking advantage of the fact that |$\ket{\Psi}_{\Lambda}$ is a ground state of the topological order [113, *PhysRevLett.115.180405, *PhysRevLett.118.110504, *PhysRevLett.118.250602, 117, *KOENIG20102707].
@@ 4.3 Wheeler-DeWitt equation and Callan-Symanzik equation from tensor networks
In the upcoming section, we will explore the application of this formalism when the bulk SymTFT contains 3D gravity. Translated into gravitational language, the no-flux condition is the statement that torsion vanishes and the curvature is constant. In the study of spin foam models, particularly those for spacetimes with positive cosmological constants, it has been shown that semi-classically, the Hamiltonian constraint corresponds to flat holonomy. In the quantum domain, this constraint is capturing re-triangulation invariance, enforced by the Pentagon identity [123, *Bonzom:2011hm]. Understanding how to achieve parallel results directly for spacetimes with negative cosmological constants, and understand 3D gravity directly in terms of $U_{q}(SL(2,\mathbb{R}))$ Levin-Wen model (especially issues regarding convergence) are interesting future directions.
@@ 4.4 Roles of tensor networks
The FRS theorem can be understood as providing a local tensor network representation of our |$\ket{\Psi}_{\Lambda}$ [36, *Fuchs:2003id, *Fuchs:2003id1, *Fuchs:2004xi, *Fjelstad:2005ua, *Fjelstad:2006aw], responsible for the topological/algebraic information of sewing rules encoded in the CFT path integral. On the other hand, our complementary construction of the state ${}_{\Lambda}\bra{\Omega}$ | [33, 34, 35, 1] provides the other half of the geometrical/complex analytical path integral information, ensuring a balanced and coherent picture of the complete structure of the system, as summarized in Fig. 2. In retrospect, the validity of this approach is underpinned by the inherent locality of 2D CFTs themselves.
Moreover, by using this exact tensor network structure and the associated RG transformation, the locality of the boundary-to-bulk map is also made explicit. This locality is the essence of the Ryu-Takayanagi (RT) formula [125] and bulk reconstruction [126, *Hamilton:2006az, 128, 129, 130, *Dong:2016eik]. In our framework, tensor networks offer a non-perturbative realization of these ideas, with locality of the map manifest at every scale. We note however that the RT formula is more than locality. In fact, the tensor network was first invoked due to the similarity between the MERA network and the RT formula[46]. At the end however, our exact tensor network turns out graph independent. One might wonder if the structure of MERA thus plays no role afterall. It turns out specific choice of graphs still plays a role in our discussion. This is because when we do a particular computation (such as the entanglement entropy of a given sub-region $A$), certain graph would lead to tremendous simplification. In the case of entanglement entropy, it turns out that if we arrange a triangulation that carries a MERA-like structure, as discussed in [1] and reviewed in Appendix D, the computation of the entanglement entropy simplifies. The location of the RT surface precisely determines the entanglement spectrum, thanks to the “quasi-perfect” property of the CFT tensors. The entanglement cut behaves almost exactly as in simple models based on perfect tensors [132], and many insights from the existing holographic tensor network literature still apply to our construction. This shows that rather than giving up on the graph, our tensor network does remember these structures. The graph independence is a way to incorporate an infinite number of graphs, which can be invoked depending on the precise computation that makes use of a particular structure. This is of course a necessary ingredient because our tensor network is an exact discrete representation of continuum field theory.
@@
Within the SymQRG procedure, a projection onto bases at various lattice scales is performed, where these bases, fixed by the symmetry of the theory, consistently reduce to the same set of labels defined by the topological symmetry (i.e. coarse-graining does not introduce new labels). This step mirrors the projection onto single-trace operators in the “quantum RG” transformations. To preserve the topological symmetry on the lattice, the transformation itself is governed by quantum $6j$ symbols. Correspondingly, the part of the tensor network responsible for the bulk theory is provided by the quantum $6j$ symbols. This approach also gives rise to a unique structure: in the bulk, it induces gauge symmetry [^24] Tensor network toy models for holography involving gauge symmetry in the bulk have been explored in [133, *Qi:2022lbd, *Dong:2023kyr, *Akers:2024ixq, 137]., while on the boundary, it preserves global symmetry. Gauge invariance essentially projects the tensor product of the boundary degrees of freedom onto an overall group singlet. Group transformation at the boundary is not part of the gauge symmetry, and they become global symmetry of the boundary degrees of freedom. This is realized in the usual WZW/Chern-Simons relation, and also in the AdS/CFT correspondence. Similarly, in the tensor network literature, geometrically higher dimensional tensor networks with gauge symmetry are used to construct global symmetric wave function in one lower dimension [138, *Singh:2012np].
@@
In this section, we demonstrate a method for reconstructing the exact bulk dual from the CFT algebra at any finite $N$. In the next section, we will see that quantum geometries naturally emerge from the CFT data, and the resulting bulk always involves a sum over these emergent quantum geometries. This framework is inherently background independent: no pre-existing geometries are assumed, and bulk operators are inherently defined within the CFT itself [143, *Goto:2016wme, *Lewkowycz:2016ukf, *Goto:2017olq], emerging alongside the geometries. The large $N$ semi-classical results can be recovered at the saddle points. However, working backward, starting from these saddle points, makes it inherently challenging to understand the emergence of spacetime once it has already emerged in the semi-classical approximation. One interesting thing we want to point out is that the sum over geometries in the bulk, in general, does not seem to include all the geometries as we usually think. The information on which geometries we need to sum in the bulk to recover a CFT is encoded in the OPE coefficients, as we will elaborate in Sec. 6.
@@ 5.3 Discrete realization of the Virasoro RG operator
These representations are special, as they are representations of the “modular double” introduced by Faddeev, which means that they are simultaneously representations of the dual quantum group $U_{\tilde{q}}(SL(2,\mathbb{R})),\tilde{q}=\text{exp}(i\pi/b^{2})$[159, 160, *Ponsot:2000mt], and it was anticipated that they would play an important role in irrational CFTs.
@@
This approach aligns with the spirit of the Regge-Ponzano model[166, *ponzano] for quantum gravity, now extended to spacetimes with a negative cosmological constant. However, we stress that analyzing bulk gravity alone is insufficient; incorporating an asymptotic boundary that realizes the $\bra{\Omega}$ | state is crucial to obtaining a finite, well-defined answer, dual to a CFT. This is the key insight of the AdS/CFT correspondence[2, 3, 4].
@@
In gravity, the insertion of this operator ensures the contractibility of the dual cycle [168]. For the Hartle-Hawking state of two-sided eternal BTZ black holes, this operator insertion applies the topological restrictions of contractibility in Euclidean signature and modifies the definition of the trace. It gives an additional $\langle\ln\hat{\omega}\rangle$ contribution in the calculation of the entanglement entropy. In the semi-classical limit, the eigenvalues $\ln(\mu(P_{\alpha}))$ are the areas, so this is the expectation value of the “quantum area operator”, which gives the correct answer we expect for the generalized entropies[169, *Wong:2022eiu, 168] [^32] The necessity of similar “defect operator” imposing a topological constraint, in the factorization map and calculation of entanglement entropy, is pointed out in JT gravity in [171], and the topological string theory in [172, *Jiang:2020cqo].. The interpretation of these line operators and their state counterparts (FZZT state[154, 155] and ZZ state[174]) as wormholes and black holes in 3D gravity are also discussed in [168].
In fact, the $T\bar{T}$ deformation preserves these Verlinde lines associated with Virasoro primaries as topological symmetries. Therefore, flows driven by $T\bar{T}$ deformations admit an interpretation as RG flow generated by these symmetry-preserving RG operators which, as we have shown, admit an interpretation as a gravitational evolution. In cases where the CFT preserves more or fewer symmetries, such as topological lines associated with more general (rational) quantum group representations in the presence of a chiral algebra that is bigger than the Virasoro algebra, then the bulk evolution could be expanded (or diminished) to make those symmetries explicit. The discussion of RG operators descending from rational TQFTs in the previous section essentially makes other rational symmetries explicit. When the RG flow does preserve these Virasoro lines, it could always be expressed in terms of $\hat{U}^{\Lambda,\Lambda^{\prime}}_{\text{RG}}(U_{q}(SL(2,\mathbb{R})))$. Although we obtained this operator by studying $|\Psi_{\text{Liouville}}\rangle$, we land on something that goes way beyond the Liouville theory. The quantum RG operator plays a universal role, since all it does is to make the symmetries preserved in the RG process explicit, the former of which determines what the latter looks like from the mathematics of representation theory. The Liouville theory, which is one example that preserves the symmetry, serves as a vehicle for us to derive the general RG operator. In other words, $|\Psi_{\text{Liouville}}\rangle$ is one solution to (4.13), but every modular invariant CFT (which we do not yet know other than Liouville theory) preserving at least the same collection of symmetries would supply an independent solution to (4.13). The take-home message is – symmetries help us identify a universal QRG operator $\hat{U}^{\Lambda,\Lambda^{\prime}}_{\text{RG}}(U_{q}(SL(2,\mathbb{R})))$ that admits an interpretation as a quantum gravitational path integral. In other words,

### v2

@@ 3 $T\bar{T}$ deformation for 2D CFTs as SymQRG
which is precisely the solution of (3.1) in terms of the Hubbard-Stratonovich transformation, and can be interpreted as coupling the original theory to 2D random geometry[81, *Cardy:2018sdv, 24].
@@ 3.3 Wheeler-DeWitt equation from $T\bar{T}$ deformation
with the integration limit switched to characterise evolution into the bulk. Then we introduce a collection of states $\bra{g_{ij}=\frac{\gamma_{ij}}{\lambda},\lambda}$ |, in analogous to the “position basis” in quantum mechanics, to be metric eigenstates in the Heisenberg picture, which plays the role of ${}_{\Lambda}\bra{\Omega^{\lambda}(\gamma_{ij})}$ | discussed in the earlier section. Then the Schrödinger picture metric eigenstates are defined at reference “time” $\lambda=\epsilon$ satisfying [62, *PhysRevD.28.2960]:
@@
The state $\bra{g_{ij}=\frac{\gamma_{ij}}{\mu},\mu}$ | corresponds to the standard Dirichlet boundary conditions imposed on finite-cutoff surfaces in the AdS/CFT correspondence. These conditions can equivalently be described as mixed boundary conditions at asymptotic infinity, linked to double-trace deformations [89, 66, 67, *Guica:2022gts]. Given that $T\bar{T}$-deformed theories have been demonstrated to preserve the Virasoro algebra in a highly nontrivial manner [67, *Guica:2022gts], it implies that these boundary conditions also lead to the same Virasoro symmetry.
@@ 4.2 SymQRG for tensor network state sum representation of 2D CFT path integrals
The observation made in [25, 26, 27, 1], motivated by [94, 95, 96] and the foundational works of FRS [28, *Fuchs:2003id, *Fuchs:2003id1, *Fuchs:2004xi, *Fjelstad:2005ua, *Fjelstad:2006aw], and also independently in [97, *Brehm:2024zun], is that 2D CFTs admit a discrete tensor network state sum representation arising from a novel perspective on cutting and gluing their path integrals. The result can be explicitly written as an overlap (2.5) in the Hilbert space of the Levin-Wen string net models.
@@
The tilde indicates a special choice of normalization that we explain in (B.3). We introduce the diagrammatic notation, where we denote $\tilde{\gamma}^{\alpha_{1}\alpha_{2}\alpha_{3}}_{I_{1}I_{2}I_{3}}(R)$ using black dashed lines, and $\tilde{C}^{\sigma_{3}\sigma_{1}\sigma_{2}}_{\alpha_{1}\alpha_{2}\alpha_{3}}$ using blue solid lines and red circles, as shown in Fig. 10. We suppress the dependence on the background metric of the 2D manifold in the chipped triangles for simplicity; see the comments in Appendix B.
For general (non-diagonal) CFTs, there is a remarkable formula first discovered by Runkel, relating the BCFT OPE coefficients to the (generalized) $6j$ symbols[99, *Runkel:1999dz, 28] as,
@@
As constructed, both states defined on the discrete lattice are projected entangled pair state (PEPS) tensor network states [^22] For readers unfamiliar with PEPS tensor networks, they can be understood as a two-dimensional generalization of the more familiar one-dimensional matrix product states(MPS) [104, *8589270]. defined on a two-dimensional lattice [106, *Verstraete:2004cf], as illustrated in Fig. 13. The dangling physical legs, shown in red, correspond to the primary labels $\alpha_{i}$. The hidden layer of the auxiliary legs represents the internal degrees of freedom, and they take values in the set of boundary conditions $\sigma_{a}$ for |$\ket{\Psi}_{\Lambda}$ and in the set of integer-valued descendant labels $I_{i}$ for ${}_{\Lambda}\bra{\Omega^{R}}$ |. For ${}_{\Lambda}\bra{\Omega^{R}}$ |, the presence of infinitely many descendants results in a bond dimension that is infinite. This infinite bond dimension is what allows for an exact representation of the fixed-point tensor network of the gapless boundary condition associated with a continuum field theory.
@@
The idea of rewriting a 2D field theory partition function as a “strange correlator”(2.5) in the topologically ordered string-net Hilbert space, to capture the topological symmetry of the theory, was inspired by the proposal in [94]. This approach was initially applied to certain 2D integrable lattice models, generalizing similar observations made for symmetry protected topological (SPT) phases in [108]. A key insight from [94] is that this formulation naturally provides an analytic real-space tensor network RG algorithm, taking advantage of the fact that |$\ket{\Psi}_{\Lambda}$ is a ground state of the topological order [109, *PhysRevLett.115.180405, *PhysRevLett.118.110504, *PhysRevLett.118.250602, 113, *KOENIG20102707].
@@ 4.3 Wheeler-DeWitt equation and Callan-Symanzik equation from tensor networks
In the upcoming section, we will explore the application of this formalism when the bulk SymTFT contains 3D gravity. Translated into gravitational language, the no-flux condition is the statement that torsion vanishes and the curvature is constant. In the study of spin foam models, particularly those for spacetimes with positive cosmological constants, it has been shown that semi-classically, the Hamiltonian constraint corresponds to flat holonomy. In the quantum domain, this constraint is capturing re-triangulation invariance, enforced by the Pentagon identity [119, *Bonzom:2011hm]. Understanding how to achieve parallel results directly for spacetimes with negative cosmological constants, and understand 3D gravity directly in terms of $U_{q}(SL(2,\mathbb{R}))$ Levin-Wen model (especially issues regarding convergence) are interesting future directions.
@@ 4.4 Roles of tensor networks
The FRS theorem can be understood as providing a local tensor network representation of our |$\ket{\Psi}_{\Lambda}$ [28, *Fuchs:2003id, *Fuchs:2003id1, *Fuchs:2004xi, *Fjelstad:2005ua, *Fjelstad:2006aw], responsible for the topological/algebraic information of sewing rules encoded in the CFT path integral. On the other hand, our complementary construction of the state ${}_{\Lambda}\bra{\Omega}$ | [25, 26, 27, 1] provides the other half of the geometrical/complex analytical path integral information, ensuring a balanced and coherent picture of the complete structure of the system, as summarized in Fig. 2. In retrospect, the validity of this approach is underpinned by the inherent locality of 2D CFTs themselves.
Moreover, by using this exact tensor network structure and the associated RG transformation, the locality of the boundary-to-bulk map is also made explicit. This locality is the essence of the Ryu-Takayanagi (RT) formula [121] and bulk reconstruction [122, *Hamilton:2006az, 124, 125, 126, *Dong:2016eik]. In our framework, tensor networks offer a non-perturbative realization of these ideas, with locality of the map manifest at every scale. We note however that the RT formula is more than locality. In fact, the tensor network was first invoked due to the similarity between the MERA network and the RT formula[38]. At the end however, our exact tensor network turns out graph independent. One might wonder if the structure of MERA thus plays no role afterall. It turns out specific choice of graphs still plays a role in our discussion. This is because when we do a particular computation (such as the entanglement entropy of a given sub-region $A$), certain graph would lead to tremendous simplification. In the case of entanglement entropy, it turns out that if we arrange a triangulation that carries a MERA-like structure, as discussed in [1] and reviewed in Appendix D, the computation of the entanglement entropy simplifies. The location of the RT surface precisely determines the entanglement spectrum, thanks to the “quasi-perfect” property of the CFT tensors. The entanglement cut behaves almost exactly as in simple models based on perfect tensors [128], and many insights from the existing holographic tensor network literature still apply to our construction. This shows that rather than giving up on the graph, our tensor network does remember these structures. The graph independence is a way to incorporate an infinite number of graphs, which can be invoked depending on the precise computation that makes use of a particular structure. This is of course a necessary ingredient because our tensor network is an exact discrete representation of continuum field theory.
@@
Within the SymQRG procedure, a projection onto bases at various lattice scales is performed, where these bases, fixed by the symmetry of the theory, consistently reduce to the same set of labels defined by the topological symmetry (i.e. coarse-graining does not introduce new labels). This step mirrors the projection onto single-trace operators in the “quantum RG” transformations. To preserve the topological symmetry on the lattice, the transformation itself is governed by quantum $6j$ symbols. Correspondingly, the part of the tensor network responsible for the bulk theory is provided by the quantum $6j$ symbols. This approach also gives rise to a unique structure: in the bulk, it induces gauge symmetry [^25] Tensor network toy models for holography involving gauge symmetry in the bulk have been explored in [129, *Qi:2022lbd, *Dong:2023kyr, *Akers:2024ixq, 133]., while on the boundary, it preserves global symmetry. Gauge invariance essentially projects the tensor product of the boundary degrees of freedom onto an overall group singlet. Group transformation at the boundary is not part of the gauge symmetry, and they become global symmetry of the boundary degrees of freedom. This is realized in the usual WZW/Chern-Simons relation, and also in the AdS/CFT correspondence. Similarly, in the tensor network literature, geometrically higher dimensional tensor networks with gauge symmetry are used to construct global symmetric wave function in one lower dimension [134, *Singh:2012np].
@@
In this section, we demonstrate a method for reconstructing the exact bulk dual from the CFT algebra at any finite $N$. In the next section, we will see that quantum geometries naturally emerge from the CFT data, and the resulting bulk always involves a sum over these emergent quantum geometries. This framework is inherently background independent: no pre-existing geometries are assumed, and bulk operators are inherently defined within the CFT itself [139, *Goto:2016wme, *Lewkowycz:2016ukf, *Goto:2017olq], emerging alongside the geometries. The large $N$ semi-classical results can be recovered at the saddle points. However, working backward, starting from these saddle points, makes it inherently challenging to understand the emergence of spacetime once it has already emerged in the semi-classical approximation. One interesting thing we want to point out is that the sum over geometries in the bulk, in general, does not seem to include all the geometries as we usually think. The information on which geometries we need to sum in the bulk to recover a CFT is encoded in the OPE coefficients, as we will elaborate in Sec. 6.
@@ 5.3 Discrete realization of the Virasoro RG operator
These representations are special, as they are representations of the “modular double” introduced by Faddeev, which means that they are simultaneously representations of the dual quantum group $U_{\tilde{q}}(SL(2,\mathbb{R})),\tilde{q}=\text{exp}(i\pi/b^{2})$[155, 156, *Ponsot:2000mt], and it was anticipated that they would play an important role in irrational CFTs.
@@
This approach aligns with the spirit of the Regge-Ponzano model[162, *ponzano] for quantum gravity, now extended to spacetimes with a negative cosmological constant. However, we stress that analyzing bulk gravity alone is insufficient; incorporating an asymptotic boundary that realizes the $\bra{\Omega}$ | state is crucial to obtaining a finite, well-defined answer, dual to a CFT. This is the key insight of the AdS/CFT correspondence[2, 3, 4].
@@
In gravity, the insertion of this operator ensures the contractibility of the dual cycle [164]. For the Hartle-Hawking state of two-sided eternal BTZ black holes, this operator insertion applies the topological restrictions of contractibility in Euclidean signature and modifies the definition of the trace. It gives an additional $\langle\ln\hat{\omega}\rangle$ contribution in the calculation of the entanglement entropy. In the semi-classical limit, the eigenvalues $\ln(\mu(P_{\alpha}))$ are the areas, so this is the expectation value of the “quantum area operator”, which gives the correct answer we expect for the generalized entropies[165, *Wong:2022eiu, 164] [^33] The necessity of similar “defect operator” imposing a topological constraint, in the factorization map and calculation of entanglement entropy, is pointed out in JT gravity in [167], and the topological string theory in [168, *Jiang:2020cqo].. The interpretation of these line operators and their state counterparts (FZZT state[150, 151] and ZZ state[170]) as wormholes and black holes in 3D gravity are also discussed in [164].
In fact, the $T\bar{T}$ deformation preserves these topological defect lines associated with Virasoro primaries as topological symmetries. Therefore, flows driven by $T\bar{T}$ deformations admit an interpretation as RG flow generated by these symmetry-preserving RG operators which, as we have shown, admit an interpretation as a gravitational evolution.
When the RG flow does preserve these Virasoro lines, it could always be expressed in terms of $\hat{U}^{\Lambda,\Lambda^{\prime}}_{\text{RG}}(U_{q}(SL(2,\mathbb{R})))$. Although we obtained this operator by studying $|\Psi_{\text{Liouville}}\rangle$, we land on something that goes way beyond the Liouville theory. The quantum RG operator plays a universal role, since all it does is to make the symmetries preserved in the RG process explicit, the former of which determines what the latter looks like from the mathematics of representation theory. The Liouville theory, which is one example that preserves the symmetry, serves as a vehicle for us to derive the general RG operator. In other words, $|\Psi_{\text{Liouville}}\rangle$ is one solution to (4.13), but every modular invariant CFT (which we do not yet know other than Liouville theory) preserving at least the same collection of symmetries would supply an independent solution to (4.13). The take-home message is – symmetries help us identify a universal QRG operator $\hat{U}^{\Lambda,\Lambda^{\prime}}_{\text{RG}}(U_{q}(SL(2,\mathbb{R})))$ that admits an interpretation as a quantum gravitational path integral. In other words,